# Matrix harmonic analysis at high temperature via the Dirichlet process

### v1

@@ 2.6. Fourier transform of $\rho^{(c)}$
Proof.
[…]
$\xi_{n}(\mathbb{R})=\{p/n:p=-n^{2},\ldots,-1,0,1,\ldots,n^{2}\}\cup\{\pm M\}.$ (2.46)
Let $\varphi$ be a continuous function with $|\varphi|\leq 1$. Then
$\mathbb{E}\,$ $\left[\varphi\left(\int\xi_{n}(x)D_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\int\varphi\left(-M\sigma_{-}+\left(\sum_{\begin{subarray}{c}p\in\mathbb{Z}\cap(-nM,nM)\end{subarray}}\frac{p}{n}\sigma_{p}\right)+M\sigma_{+}\right){\mathrm{Dir}_{(c\vec{a})}(\mathrm{d}\vec{\sigma})}$ (2.47)
$a_{p}=\rho^{[M]}(\xi_{n}^{-1}(\{p/n\})),\quad p\in\mathbb{Z}\cap(-nM,nM)$ (2.48)
$a_{-}=\rho^{[M]}(\{-M\}),$ (2.49)
$a_{+}=\rho^{[M]}(\{M\}).$ (2.50)
Since $\xi_{n}^{-1}(\{p/n\})$ in contained in (M,M)(-M,M), we have
$a_{p}=\rho(\xi_{n}^{-1}(\{p/n\})),\quad p\in\mathbb{Z}\cap(-nM,nM)$ (2.51)
$a_{-}=\rho((-\infty,-M]),$ (2.52)
$a_{+}=\rho([M,+\infty)).$ (2.53)
Thus by the definition of the Dirichlet process, one has
$\mathbb{E}\,$ $\left[\varphi\left(\int\xi_{n}(x)P_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int\left(-M\mathbbm{1}_{x\leq-M}+\xi_{n}(x)\mathbbm{1}_{x\in(-M,M)}+M\mathbbm{1}_{x\geq-M}\right)D_{c\rho}(\mathrm{d}x)\right)\right]$ (2.54)
We then take the $n\to\infty$ limit on both sides. On the left hand side, because of the continuity of $\varphi$ and the boundedness of $\xi_{n}(x)$ by $|x|$, one can push the limit inside both integrals and obtain
$\lim_{n\to+\infty}\mathbb{E}\,\left[\varphi\left(\int\xi_{n}(x)D_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}}(\mathrm{d}x)\right)\right].$ (2.55)
$\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}}(\mathrm{d}x)\right)\right]=\mathbb{E}\,\left[\varphi\left(\int\left(-M\mathbbm{1}_{x\leq-M}+x\mathbbm{1}_{x\in(-M,M)}+M\mathbbm{1}_{x\geq-M}\right)D_{c\rho}(\mathrm{d}x)\right)\right].$ (2.56)
Taking $M\to+\infty$ limit on both sides gives
$\lim_{M\to\infty}\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho^{[M]}(\mathrm{d}x)}\right)\right]=\mathbb{E}\,\left[\varphi\left(\int xD_{c\rho(\mathrm{d}x)}\right)\right].$ (2.57)
This proofs that $(\rho^{[M]})^{(c)}\to\rho^{(c)}$ weakly. ∎
Lemma 2.11.
[…]
$e^{ux}={u\Gamma(c)}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\frac{e^{uz}}{(u(z-x))^{c}}$ (2.58)
where
For some large enough $M>x$, the contour $\mathcal{C}$ is the Hankel loop starts and ends at -$\infty$, enclosing $(-\infty,x]$ oriented counter-clockwise. A typical choice is that, it goes horizontally from $\infty-iM$ to MiMM-iM, then vertically to $M+iM$, and then horizontally to $\infty+iM$.
Proof.
$\frac{1}{\Gamma(c)}=\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\frac{e^{z}}{z^{c}}.$ (2.60)
$\begin{split}\left|\int_{C_{R}}\frac{e^{uz}}{(u(z-x)^{c})}\right|&\leq\int_{\pi/2}^{\pi}\mathrm{d}\theta\frac{Re^{\mathrm{Re}\,(uRe^{i\theta})}}{|u(Re^{i\theta}-x)|^{c}}\\ &\leq\frac{\pi}{2}\frac{R}{|u(R-|x|)|^{c}}\sup_{\theta\in(\pi/2,\pi)}\exp(R(\mathrm{Re}\,u\cos\theta-\mathrm{Im}\,u\sin\theta)).\end{split}$ (2.61)
In the case $\mathrm{Re}\,u<0$, one can see that after the change of variable, the contour is deformed such that it starts and ends at the first (or fourth) quadrant if $\mathrm{Im}\,u\geq 0$ (or $\mathrm{Im}\,u<0$). A similar bound to (2.61) can be performed to see that applying a Cauchy theorem the contour can be deformed such that it encloses the branch point u(zx)u(z-x) with a branch cut going to $\infty$. We have
$\frac{1}{\Gamma(c)}=u\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\frac{e^{u(z-x)}}{(u(z-x))^{c}}.$ (2.62)
A simplification gives us the result.
$\frac{1}{\Gamma(c)}=\frac{1}{2\pi i}\int_{-i\mathcal{C}}it\mathrm{d}z\,\frac{e^{it(z-x)}}{(it(z-x))^{c}}.$ (2.63)
The contour $i\mathcal{C}$ encloses the branch cut $i(-\infty,x]$.
Lemma 2.12.
For $t\neq 0$ and $\varepsilon>0$ small enough,
$\int_{I_{t}}\exp(R(\varepsilon\cos\theta-t\sin\theta))\mathrm{d}\theta=O\left(\frac{e^{\varepsilon R}}{tR}\right)$ (2.64)
where $I_{t}=[\pi/2,\pi]$ if $t>0$, and $I_{t}=[\pi,3\pi/2]$ if $t<0$.
Proof.
$\begin{split}&\theta_{\max}=\arctan(-t/\varepsilon)\in\left(-\frac{\pi}{2},0\right),\quad\theta_{\min}=\arctan(-t/\varepsilon)+\pi\in\left(\frac{\pi}{2},\pi\right).\end{split}$ (2.65)
$\int_{a}^{b}\exp(R(\varepsilon\cos\theta-t\sin\theta))\mathrm{d}\theta=O\left(\frac{e^{R\phi(\pi)}}{R|\phi^{\prime}(\pi)|}\right).$ (2.66)
We also have
$\phi(\pi)=\varepsilon,\quad\phi^{\prime}(\pi)=-t\neq 0.$ (2.67)
Assembling the above gives the result.
With the previous two lemmas, one can state the following theorem which gives a contour integral expression for the Fourier transform (or the Laplace-transform-type integral in general) of $\rho^{(c)}$. This theorem may be formally derived from [41, Theorem 7] using a change of variable, but in their paper the convergence of integrals have not been taken care of. Here we provide an detailed proof similar to the one used in [41], taking care of all issues of convergence.
Theorem 2.13.
[…]
$\int_{\mathbb{R}}e^{x\mathrm{Re}\,u}\,\rho(\mathrm{d}x)<+\infty$ (2.68)
Then
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{xu}=\lim_{M\to+\infty}\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int\log(z-s)\rho^{[M]}(\mathrm{d}s)\right)$ (2.69)
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{xu}=\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int\log(z-s)\rho(\mathrm{d}s)\right).$ (2.70)
Proof.
Step I: Compactly supported $\rho$. By Theorem 2.7, the left side of (2.70) is finite. By Lemma 2.11 we have
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}=\int_{-M}^{M}\rho^{(c)}(\mathrm{d}x)\,{u\Gamma(c)}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\frac{e^{uz}}{(u(z-x))^{c}}$ (2.71)
$\begin{split}\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}&=y\Gamma(c)\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,e^{uz}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{(u(z-x))^{c}}\\ &=u\Gamma(c)\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(u(z-s))\rho(\mathrm{d}s)\right),\end{split}$ (2.72)
$\log(u(z-s))=\log u+\log(z-s)+2m\pi i$ (2.73)
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{ux}$ $=\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right).$ (2.74)
Step II: $u=it$ with $t>0$. In the special case where $\rho$ is compactly supported and $u=it$ with $t>0$, we need to deform the contour $\mathrm{sgn}(u)\mathcal{C}$ to $\mathbb{R}-i\delta$, for every $\delta>0$.
Denote the arc going from R+MiMR+M-iM to M+i(RM)M+i(R-M) centering at MiMM-iM by $\mathcal{C}_{R}^{+}$, and the arc going from M+i(RM)-M+i(R-M) to R+MiMR+M-iM centering at MiM-M-iM by $\mathcal{C}_{R}^{-}$; they both have radius $R$. Then we have the inequality
$\begin{split}\left|\int_{\mathcal{C}_{R}^{+}}\mathrm{d}z\,e^{itz}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{(it(z-x))^{c}}\right|&=\left|\int_{0}^{\pi/2}Rie^{i\theta}\mathrm{d}\theta\,e^{it(M-iM)}e^{itRe^{i\theta}}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(M-iM+Re^{i\theta}-x)^{c}}\right|\\ &\leq\int_{0}^{\pi/2}\mathrm{d}\theta\,Re^{tM-tR\sin\theta}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(R-3M)^{c}}=\frac{e^{tM}}{t^{c}(R-3M)^{c}}O(1)\to 0\end{split}$ (2.75)
as $R\to+\infty$, where we use the asymptotics
$\int_{0}^{\pi/2}\mathrm{d}\theta\,e^{-R\sin\theta}=O(R^{-1}),\text{ as }R\to+\infty,$ (2.76)
which can be deduced via a Laplace’s method (c.f. Lemma 2.12). Similarly for $\mathcal{C}_{R}^{-}$ we have
$\begin{split}\left|\int_{\mathcal{C}_{R}^{-}}\mathrm{d}z\,e^{itz}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{(it(z-x))^{c}}\right|&=\left|\int_{0}^{\pi/2}Rie^{i\theta}\mathrm{d}\theta\,e^{it(-M-iM)}e^{itRe^{i(\theta+\pi/2)}}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(-M-iM+Re^{i(\theta+\pi/2)}-x)^{c}}\right|\\ &\leq\int_{0}^{\pi/2}\mathrm{d}\theta\,Re^{tM-tR\cos\theta}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(R-3M)^{c}}=\frac{e^{tM}}{t^{c}(R-3M)^{c}}O(1)\to 0\end{split}$ (2.77)
$\begin{split}\left|\int_{\gamma_{\pm R}}\mathrm{d}z\,e^{itz}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{(it(z-x))^{c}}\right|&=\left|\int_{M}^{\delta}(-i)\mathrm{d}y\,e^{it(\pm R-iy)}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(\pm R-iy-x)^{c}}\right|\\ &\leq\int_{M}^{\delta}\mathrm{d}y\,e^{ty}\int_{-M}^{M}\frac{\rho^{(c)}(\mathrm{d}x)}{t^{c}(R-2M)^{c}}\to 0.\end{split}$ (2.78)
Thus one has
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{itx}=\frac{\Gamma(c)}{(it)^{c-1}}\frac{1}{2\pi i}\int_{\mathbb{R}-i\delta}\mathrm{d}z\,\exp\left(itz-c\int\log(it(z-s))\rho(\mathrm{d}s)\right).$ (2.79)
which gives the final result for $t>0$.
Step IV: Non-compact suppport. For the case where $\rho$ does not have a compact support, one makes use of the measure $\rho^{[M]}$ and the weak convergence $\rho^{[M]}\to\rho$ (by Lemma 2.10). Noting that $\rho^{[M]}$ has a compact support, and thus by the Lebesgue dominated convergence theorem, we have the final result.
Step V: Special case. In addition, for the case $u=it$, $t\neq 0$ and $c>1$,
$\begin{split}\left|\int_{\mathbb{R}-i\delta\mathrm{sgn}(t)}\mathrm{d}z\,e^{itz}\int_{\mathbb{R}}\frac{(\rho^{[M]})^{(c)}(\mathrm{d}x)}{(z-x)^{c}}\right|&\leq e^{\delta\mathrm{sgn}(t)}\int_{\mathbb{R}}\mathrm{d}y\,\int_{\mathbb{R}}\frac{(\rho^{[M]})^{(c)}(\mathrm{d}x)}{((y-x)^{2}+\delta^{2})^{c/2}}.\end{split}$ (2.80)
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{itx}=\frac{\Gamma(c)}{(it)^{c-1}}\frac{1}{2\pi i}\int_{\mathbb{R}-i\delta\,\mathrm{sgn}(t)}\mathrm{d}z\,e^{itz}\lim_{M\to+\infty}\int_{\mathbb{R}}\frac{(\rho^{[M]})^{(c)}(\mathrm{d}x)}{(z-x)^{c}}.$ (2.81)
By Lemma 2.10, the weak convergence $(\rho^{[M]})^{(c)}\to\rho^{(c)}$ finalises the proof. ∎
2.7. Mellin transform of $\rho^{(c)}$
Similar to the Fourier transform, the Mellin transform of $\rho^{(c)}$ can also be rewritten into a contour integral formula.
Lemma 2.14.
[…]
@@ 2.7. Mellin transform of $\rho^{(c)}$
as $R\to\infty$. Therefore the contour can be deformed to the Hankel loop, for example the one given in Lemma 2.11 (provided $M>\log x$). ∎
@@
$\int_{0}^{+\infty}x^{\mathrm{Re}\,u}\,\rho(\mathrm{d}x)<+\infty$ (2.87)
@@
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)\,x^{u}=\lim_{M\to+\infty}\frac{\Gamma(c)\Gamma(u+1)}{\Gamma(u+c)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int\log(1-se^{-z})\rho^{[M]}(\mathrm{d}s)\right)$ (2.88)
@@
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)\,x^{y}=\frac{\Gamma(c)\Gamma(u+1)}{\Gamma(u+c)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int\log(1-se^{-z})\rho(\mathrm{d}s)\right).$ (2.89)
@@
for some Hankel loop $\mathcal{C}$ enclosing $(-\infty,\log M]$. By the exponential decay of $e^{uz}$ along $\mathcal{C}$ (because $\mathrm{Re}\,u>0$) and that $|1-xe^{-z}|$ is bounded from below, the Fubini theorem can be applied to interchange the two integrals. One also recalls that by (2.23)
@@
and thus we obtain (2.89) for compactly supported $\rho$. Equation (2.88) can be again obtained by Lemma 2.10 and the Lebesgue dominated convergence theorem.
For general $\rho$, we have the weak convergence $(\rho^{[M]})^{(c)}\to\rho^{(c)}$ by Lemma 2.10, which finalises the proof. ∎
@@ 3. Proof of the main theorems
3.1. Relation with symmetric functions
@@
A consequence of this, shown in [14, Thm 5.1 and Thm 5.2], is an integral representation of the multivariate Bessel function. The following theorem extends the two theorems in [14] to the case $\mathrm{Re}\,u=0$.
Theorem 3.1 (for $\mathrm{Re}\,u\neq 0$ see Thm 5.1 and Thm 5.2 in [14]).
[…]
$\mathcal{B}_{\vec{a}}(u;N,\theta)=\frac{\Gamma(\theta N)}{u^{N\theta-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,e^{uz}\prod_{j=1}^{N}(z-a_{j})^{-\theta}.$ (3.2)
@@
holds as $R\to\infty$, where the final bound is obtained from applying Lemma 2.12. A similar inequality holds for the contour $\mathcal{C}_{2}$ as
@@
$\mathcal{F}_{\log\vec{a}}(u;N,\theta)=\frac{\Gamma(\theta N)\Gamma(u+1)}{\Gamma(u+\theta N)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,e^{uz}\prod_{j=1}^{N}(1-a_{j}e^{-z})^{-\theta}.$ (3.9)
@@
$\frac{\Gamma_{e^{-\varepsilon}}(\varepsilon^{-1}z-(\lfloor\varepsilon^{-1}\log a_{j}\rfloor-\theta j+\theta))}{\Gamma_{e^{-\varepsilon}}(\varepsilon^{-1}z-(\lfloor\varepsilon^{-1}\log a_{j}\rfloor-\theta j))}=(1-e^{-\varepsilon})^{\theta}\frac{(e^{-(z-\varepsilon\lfloor\varepsilon^{-1}\log a_{j}\rfloor)}e^{-\varepsilon\theta(1-j)};e^{-\varepsilon})_{\infty}}{(e^{-(z-\varepsilon\lfloor\varepsilon^{-1}\log a_{j}\rfloor)}e^{-\varepsilon\theta(-j)};e^{-\varepsilon})_{\infty}}$ (3.11)
@@
$\begin{split}\mathcal{F}_{\log\vec{a}}(u;N,\theta)=&\lim_{\varepsilon\to 0^{+}}\frac{-\varepsilon}{e^{-\varepsilon}-1}(1-e^{-\varepsilon})^{\theta N}\frac{(e^{-\varepsilon(u+\theta N)};e^{-\varepsilon})_{\infty}}{(e^{-\varepsilon(y+1)};e^{\varepsilon})_{\infty}}\frac{\Gamma_{e^{-\varepsilon}}(\theta N)}{2\pi i}\times\\ &\frac{1}{\varepsilon}\int_{\mathcal{C}}\mathrm{d}z\,e^{-uz}\prod_{j=1}^{N}\frac{(e^{-x_{j}(\varepsilon)}e^{-\varepsilon\theta(1-j)};e^{-\varepsilon})_{\infty}}{(e^{-x_{j}(\varepsilon)}e^{-\varepsilon\theta(-j)};e^{-\varepsilon})_{\infty}}\end{split}$ (3.12)
@@
$\lim_{\varepsilon\to 0^{+}}(1-e^{-\varepsilon})^{\theta N-1}\frac{(e^{-\varepsilon(u+\theta N)};e^{-\varepsilon})_{\infty}}{(e^{-\varepsilon(u+1)};e^{\varepsilon})_{\infty}}=\frac{\Gamma(u+1)}{\Gamma(u+\theta N)}$ (3.13)
@@
Combining Theorems 3.1, 3.2, 2.13, and 2.15, we obtain the following corollary.
@@
$\mathcal{F}_{\log\vec{a}}(u;N,\theta)=\int_{0}^{+\infty}x^{u}\rho_{N}^{(c_{N})}(\mathrm{d}x)$ (3.16)

### v3

@@ 2.5. Mellin transform of $\rho^{(c)}$
For all $c>0$, $u\in\mathbb{C}\setminus\{0\}$ with $\mathrm{Re}\,u>0$, $x>0$ and $t\in\mathbb{R}\setminus\{0\}$, we have
@@
as $R\to\infty$. Therefore the contour can be deformed to the Hankel loop, for example the one given in Lemma A.3 (provided $M>\log x$). ∎
@@
$\int_{0}^{+\infty}x^{\mathrm{Re}\,u}\,\rho(\mathrm{d}x)<+\infty,$ (2.36)
@@
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)\,x^{u}=\lim_{M\to+\infty}\frac{\Gamma(c)\Gamma(u+1)}{\Gamma(u+c)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(1-se^{-z})\rho^{[M]}(\mathrm{d}s)\right)$ (2.37)
@@
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)\,x^{y}=\frac{\Gamma(c)\Gamma(u+1)}{\Gamma(u+c)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(1-se^{-z})\rho(\mathrm{d}s)\right).$ (2.38)
Here $\mathcal{C}$ is the Hankel loop enclosing $(-\infty,\log M]$.
@@
for some Hankel loop $\mathcal{C}$ enclosing $(-\infty,\log M]$. By the exponential decay of $e^{uz}$ along $\mathcal{C}$ (because $\mathrm{Re}\,u>0$) and that $|1-xe^{-z}|$ is bounded from below, the Fubini theorem can be applied to interchange the two integrals. One also recalls that by (A.16)
@@
and thus we obtain (2.38) for compactly supported $\rho$. Equation (2.37) can be again obtained by Lemma A.2 and the Lebesgue dominated convergence theorem.
For general $\rho$, we have the weak convergence $(\rho^{[M]})^{(c)}\to\rho^{(c)}$ by Lemma A.2, which finalises the proof. ∎
@@ 3. Proof of the main theorems
3.1. Integral representations for symmetric functions
@@
A consequence of this, shown in [19, Thm 5.1 and Thm 5.2], is an integral representation of the multivariate Bessel function. The following theorem extends the two theorems in [19] to the case $\mathrm{Re}\,u=0$.
Proposition 3.1 (for $\mathrm{Re}\,u\neq 0$ see Thm 5.1 and Thm 5.2 in [19]).
Let $\theta>0$, $N\in\mathbb{N}$, $a_{1}\leq\ldots\leq a_{N}$ in $\mathbb{R}$ and $u\in\mathbb{C}\setminus\{0\}$. Then
$\mathcal{B}_{\vec{a}}(u;N,\theta)=\frac{\Gamma(\theta N)}{u^{N\theta-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,e^{uz}\prod_{j=1}^{N}(z-a_{j})^{-\theta},$ (3.2)
@@
holds as $R\to\infty$, where the final bound is obtained from applying Lemma A.4. A similar inequality holds for the contour $\mathcal{C}_{2}$ as
@@
$\mathcal{F}_{\log\vec{a}}(u;N,\theta)=\frac{\Gamma(\theta N)\Gamma(u+1)}{\Gamma(u+\theta N)}\frac{1}{2\pi i}\int_{\mathcal{C}}\mathrm{d}z\,e^{uz}\prod_{j=1}^{N}(1-a_{j}e^{-z})^{-\theta},$ (3.9)
@@
$\frac{\Gamma_{e^{-\varepsilon}}(\varepsilon^{-1}z-(\lfloor\varepsilon^{-1}\log a_{j}\rfloor-\theta j+\theta))}{\Gamma_{e^{-\varepsilon}}(\varepsilon^{-1}z-(\lfloor\varepsilon^{-1}\log a_{j}\rfloor-\theta j))}=(1-e^{-\varepsilon})^{\theta}\frac{(e^{-(z-\varepsilon\lfloor\varepsilon^{-1}\log a_{j}\rfloor)}e^{-\varepsilon\theta(1-j)};e^{-\varepsilon})_{\infty}}{(e^{-(z-\varepsilon\lfloor\varepsilon^{-1}\log a_{j}\rfloor)}e^{-\varepsilon\theta(-j)};e^{-\varepsilon})_{\infty}}.$ (3.11)
@@
$\begin{split}\mathcal{F}_{\log\vec{a}}(u;N,\theta)=&\lim_{\varepsilon\to 0^{+}}\frac{-\varepsilon}{e^{-\varepsilon}-1}(1-e^{-\varepsilon})^{\theta N}\frac{(e^{-\varepsilon(u+\theta N)};e^{-\varepsilon})_{\infty}}{(e^{-\varepsilon(y+1)};e^{\varepsilon})_{\infty}}\frac{\Gamma_{e^{-\varepsilon}}(\theta N)}{2\pi i}\times\\ &\frac{1}{\varepsilon}\int_{\mathcal{C}}\mathrm{d}z\,e^{-uz}\prod_{j=1}^{N}\frac{(e^{-x_{j}(\varepsilon)}e^{-\varepsilon\theta(1-j)};e^{-\varepsilon})_{\infty}}{(e^{-x_{j}(\varepsilon)}e^{-\varepsilon\theta(-j)};e^{-\varepsilon})_{\infty}}.\end{split}$ (3.12)
@@
$\lim_{\varepsilon\to 0^{+}}(1-e^{-\varepsilon})^{\theta N-1}\frac{(e^{-\varepsilon(u+\theta N)};e^{-\varepsilon})_{\infty}}{(e^{-\varepsilon(u+1)};e^{\varepsilon})_{\infty}}=\frac{\Gamma(u+1)}{\Gamma(u+\theta N)}.$ (3.13)
@@
Combining Propositions 3.1, 3.2, 2.9, and 2.11, we obtain the following corollary.
@@
$\mathcal{F}_{\log\vec{a}}(u;N,\theta)=\int_{0}^{+\infty}x^{u}\rho_{N}^{(c_{N})}(\mathrm{d}x).$ (3.16)